OpenStudy (anonymous):

plzzzz need help ODEs \(\frac{dy}{dx}=\frac{xy^2-cos x sin x}{y(1-x^22)}\) \(y(0)=2\)

OpenStudy (anonymous):

check if his is correct \(y dy -x^2dy = xy^2dx -cos x sinx dx\)

OpenStudy (anonymous):

\(\frac{1}{2}y^2 -x^2y=\frac {1}{2}x^2 y^2-\int\cos x \sin x \dx \)

OpenStudy (anonymous):

?? i dont know if im on the right direction

ganeshie8 (ganeshie8):

nope, y is a function of x u cannot integrate \(\large \int xy^2 dx\)

ganeshie8 (ganeshie8):

\(\large \frac{dy}{dx}=\frac{xy^2-cos x sin x}{y(1-x^2)}\)

ganeshie8 (ganeshie8):

like this right ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

\(\large y' =\frac{xy^2-cos x sin x}{y(1-x^2)}\) multiply both sides by \(y\) \(\large yy' =\frac{xy^2-cos x sin x}{1-x^2}\)

OpenStudy (anonymous):

isnt y'=dx/dy ?

OpenStudy (anonymous):

sorry im not really good :'(

ganeshie8 (ganeshie8):

substitute \(\large v = y^2\) that gives, \(\large v' = 2yy'\)

ganeshie8 (ganeshie8):

the equation becomes : \(\large \frac{v'}{2} =\frac{xv-cos x sin x}{1-x^2}\) \(\large v' =\frac{2xv-\sin (2x)}{1-x^2}\)

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good, knw how to solve a "linear equation" using IF method ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

show mw if pspl ill understand quick

ganeshie8 (ganeshie8):

this is a bernouli equation

ganeshie8 (ganeshie8):

after substitution it changes to a linear equation

ganeshie8 (ganeshie8):

you should know how solve "linear equations" and the Intgeration Factor method, before starting bernouli

OpenStudy (anonymous):

well , im sure i learned them but my basecs are rusty ( been long time ) so if u show me i might remember

ganeshie8 (ganeshie8):

\(\large y' =\frac{xy^2-cos x sin x}{y(1-x^2)}\) multiply both sides by \(y\) \(\large yy' =\frac{xy^2-cos x sin x}{1-x^2}\) substitute \(\large v = y^2\) \(\large v' = 2yy'\) the equation becomes : \(\large \frac{v'}{2} =\frac{xv-cos x sin x}{1-x^2} \) \(\large v' =\frac{2xv-\sin (2x)}{1-x^2} \) \(\large v' + \frac{2x}{1-x^2}v = \frac{\sin(2x)}{x^2-1} \)

ganeshie8 (ganeshie8):

thats what we have so far^

OpenStudy (anonymous):

i with u so far

ganeshie8 (ganeshie8):

\(\large y' =\frac{xy^2-cos x sin x}{y(1-x^2)}\) multiply both sides by \(y\) \(\large yy' =\frac{xy^2-cos x sin x}{1-x^2}\) substitute \(\large v = y^2\) \(\large v' = 2yy'\) the equation becomes : \(\large \frac{v'}{2} =\frac{xv-cos x sin x}{1-x^2} \) \(\large v' =\frac{2xv-\sin (2x)}{1-x^2} \) \(\large v' + \frac{2x}{x^2-1}v = \frac{\sin(2x)}{x^2-1} \)

OpenStudy (anonymous):

im*

ganeshie8 (ganeshie8):

just corrected a typo in the last line

ganeshie8 (ganeshie8):

Multiply the IF \((x^2-1)\) through out the equaiton

ganeshie8 (ganeshie8):

the equation becomes : \(\large (x^2-1)v' + 2xv = \sin(2x) \)

ganeshie8 (ganeshie8):

Using product rule, left hand side can be compressed to : \(\large ((x^2-1)v)' \)

ganeshie8 (ganeshie8):

\(\large (x^2-1)v' + 2xv = \sin(2x) \) becomes : \(\large ((x^2-1)v)' = \sin(2x) \)

ganeshie8 (ganeshie8):

Now is the good time to integrate both sides

OpenStudy (anonymous):

how ?

ganeshie8 (ganeshie8):

integrate both sides

ganeshie8 (ganeshie8):

\(\large \int ((x^2-1)v)'~ dx = \int \sin(2x)~dx \)

ganeshie8 (ganeshie8):

By FTC, \(\large \int f' dx = f\), so left hand sides gives u back \((x^2-1)v\)

ganeshie8 (ganeshie8):

\(\large \int ((x^2-1)v)'~ dx = \int \sin(2x)~dx \) \(\large (x^2-1)v = \int \sin(2x)~dx \)

ganeshie8 (ganeshie8):

you can integrate the right hand side

OpenStudy (anonymous):

the only thing that i can do by myself :P thank you for your time I really appreciate it :D

ganeshie8 (ganeshie8):

np :) dont forget substituting the v's back to y^2

OpenStudy (anonymous):

okeyy i wont

OpenStudy (anonymous):

\(y=\sqrt {\frac{cos 2x}{2 x^2-2 }}+c\) \(y(0)=\sqrt {\frac{cos 0}{ -2 }}+c=2\rightarrow c=2-\sqrt{\frac{1}{2}}\)

ganeshie8 (ganeshie8):

\(\large \large (x^2-1)v = \int \sin(2x)~dx\) \(\large \large (x^2-1)y^2 = -\frac{\cos (2x)}{2} + C\)

ganeshie8 (ganeshie8):

plugin (0, 2) : \(\large C = \frac{-7}{2}\)

ganeshie8 (ganeshie8):

\(\large \large (x^2-1)y^2 = -\frac{\cos (2x)}{2} - \frac{7}{2}\)

OpenStudy (anonymous):

im really stink xD thanks again

ganeshie8 (ganeshie8):

you can solve y if u want